OpenStudy (anonymous):

Solve the given system, using the substitution method. y = 3x – 7 6x – 2y = 12

OpenStudy (anonymous):

A. (14, 12) B. (12, 14) C. There is no solution. D. There are an infinite number of solutions.

OpenStudy (anonymous):

no solutions

OpenStudy (luigi0210):

Well sub in the first into the second: \[\LARGE 6x-2(3x-7)=12\] Now distribute: \[\LARGE 6x-6x+14=12\] Simplify: \[\LARGE 14 \neq 12\]

OpenStudy (anonymous):

The lines are parallel so they do not intersect. Meaning no solutions.

OpenStudy (luigi0210):

So yea, no solutions. Unless in the math world, 14 somehow does equal 12.

OpenStudy (anonymous):

ok Thank you @Luigi0210 and @fritholly1

OpenStudy (anonymous):

Solve the system using the substitution method. Show all your steps. 1. x+2y=6 3x-2y=2 3. 4x+y=7 2x+5y=-1 5. 3x+2y=-2 6x-y=6 Solve the system using the elimination method. Show all your steps. 7. -3x+3y=3 3x+y=9 9. -5x+12y=20 x-2y=-6 11. 3x+2y=1 4x+ 6y= 7 Solve the system using any algebraic method. Show all your steps. 13. 5x+ 7y= -2 2x-7y=9 14. x+3y=1 3x+7y=1 15. 4x+6y=8 2x+3y= 3 16. 8x-5y= -17 -2x+y=6 17. 3x-8y=0 -2x+5y=-2 18. 4x-6y=2 5x+3y= 1 19. 2x-5y=3 -4x+10y= -6 20. 8x+3y=10 -6x + y = -12 21. 5x+ 4y= -18 2x+3y=-24 22. CDs and Cassettes From 1990 to 1998, the manufacturer's shipments for audio cassettes A(in millions) and compact discs C(in millions) can be modeled by the equations A= -31.8t + 322 Audio cassette shipments C= 42.8t+110 Compact disc shipments where t is the number of years since 1990. During what year did the number of compact discs shipped surpass the number of audio cassettes shipped? Solve the system using any algebraic method. Show your steps. 30. 3x+7y= -1 2x +3y= 6 32. 3x-y= -2 5x+2y= 15 34. 2x + 3y= -6 3x- 4y= 25 36. 4x- 3y= 8 -8x +6y= 16 38. 7x+5y= -12 3x-4y=1 Solve the system. Then classify the system as consistent and independent, consistent and dependent, or inconsistent. 4, 4x+ 8y=8 x+ 2y=6 5. -5x+3y= -5 y= 5/3x + 1 6. x-2y=2 2x-y= -5 Solve the system using the substitution method. Show all steps. 7. 3x-y= -4 x+ 3y= -28 8. x+ 5y=1 -3x +4y= 16 9. 6x +y= -6 4x+3y= 17 Solve the system using the elmination method. Show all steps. 10. 2x- 3y= -1 2x+3y= -19 11. 3x-2y= 10 -6x +4y= -20 12. 2x +3y= 17 5x+8y= 20

OpenStudy (anonymous):

Can you answer a few more questions I have?

OpenStudy (luigi0210):

Open a new question, and we'll be there.